what is the dimension to the word problem A rectangle has a perimeter of 18 cm. Its length is 5 cm greater than its width. Find the dimension
Answer by colliefan(242)

If the width is w, the length can be written as w+5. The perimeter is 18. It is the sum of 2 lengths and 2 widths so:
2w + 2(w+5) = 18
Then simplify to solve for w. What you do to one side, you must do to the other.
2w + 2(w+5) = 18
2w+2w+10=18
4w+10-10=18-10
4w+0=8
4w=8
1/4*4w=8*1/2
1/4*4w=8*1/4
1*w=2
w=2

Then, the length must equal w+5 or 2+5 or 7
Check to see that the perimeter of a rectangle with 2 sides of 2 cm and 2 sides of 7 cm equals 18.